OpenStudy (anonymous):

Help me solve this question below: Find the fraction that is equivalent to 3/4 and : a) the difference between the denominator and the numerator is 7 b) the product of the denominator and the numerator is 300

OpenStudy (anonymous):

Let the numerator be x then denominator =x+7 So, x/(x+7) =3/4

OpenStudy (anonymous):

solve for x and find x/(x+7)

OpenStudy (anonymous):

can u?

OpenStudy (anonymous):

yes i can with a but .......how about b

OpenStudy (anonymous):

Let numerator be 3x then denominator =4x Now, 3x*4x=300 SOLVE FOR x and find 3x/4x

OpenStudy (anonymous):

NO